September 2018 water damage

From LNF Wiki
Jump to navigation Jump to search

This page has been created to consolidate updates following the September 27th flood that occurred in EECS 1450 (submicron room) and 1440D/E (thin films and plasma bays). Note that the rest of the LNF spaces were not affected. Please see below for updated status.

Facility status

Update as of mid day, Wednesday October 3rd: the cleaning of EECS 1440 D/E has been completed and these areas of the clean room were re-opened to users.
EECS 1450 (submicron room) has been cleaned, particle count tested and is open to users.

Equipment status

Most equipment is now back to normal status. Please check the LNF scheduler for up to date information.

Equipment Status Alternatives Comments
GCA AS200 AutoStep Back up - 10/8/18 Global alignment calibration is currently less stable due to environmental chamber,

µDFAS however is working well. (create a Helpdesk ticket if you need assistance)

Oxford OpAl ALD Back up - 10/9/18 Test run shows process is running as before. Tool did cool rapidly so the chamber did flake out. It was cleaned up but particles may still be high.

Alternative solutions and back up

Contact us for any urgent need - please create a Helpdesk ticket on any tool that you need and provide as much process information as possible.